OpenStudy (surryyy):

Use accurate and appropriate terminology to describe the mathematical ideas presented

OpenStudy (king.void.):

I'm not good at math sorry..

OpenStudy (texaschic101):

where are the mathematical ideas ??

OpenStudy (surryyy):

i think its about Additive and Multiplicative Inverses

OpenStudy (texaschic101):

well...the multiplication inverses is when you multiply a certain number by its reciprocal and get the answer of 1. example : 8 * 1/8 = 1 , 2/1 * 1/2 = 1 , -3/4 * -4/3 = 1 the additive inverse is when you add a number by its opposite to get the answer of 0. example : -5 + 5 = 0, 2 + (-2) = 0